# What is “Typological Diversity” in NLP?

### v3

@@ 1 Introduction
Current work increasingly evaluates models on multiple languages, but because of resource constraints, it is not realistic to test a model on the thousands of languages in the world. In order to still ensure a degree of generalizability, previous work recognizes the importance of diverse language sampling. Ponti et al. (2020) suggest that merely evaluating on a small set of similar languages is an unreliable method for estimating a multilingual model’s performance, since such evaluations lack robustness to previously unseen typological properties, for instance. In an attempt to address these sampling issues, NLP research increasingly claims to rely on linguistic typology (see Figure 1). Evaluating on a set of so-called ‘typologically diverse’ languages implies robustness over typological variations, implicitly claiming broad generalizability across languages. Importantly however, there is a fundamental lack of common understanding of exactly what a claim of typological diversity entails.
In this work, we systematically investigate how the term ‘typological diversity’ is used in NLP research by surveying the literature for past and current practices. We find that there is no set definition of ‘typological diversity’ in NLP: justifications for claims range from (i) none whatsoever, to (ii) using language phylogeny as a proxy for typology, to (iii) basing diversity on a more or less well-founded subset of available typological features. As a consequence, typological language similarity and typological feature value inclusion vary considerably across papers (§5). Furthermore, we demonstrate why this is problematic, by showing how a skewed selection of evaluation languages can lead to overstated multilingual performance and poor generalizability (§6). We recommend future approaches that use the term ‘typologically diverse’ to document their language samples, and to add a measure of typological diversity, such as the mean pairwise language distance (§5.1) or typological feature inclusion (§5.2). Better-documented language sampling is crucial for the development of multilingual NLP, because it provides insights into commonly overlooked linguistic difficulties of multilingual language modelling. We provide software to verify and visualize this, linked below the abstract.
@@ 2 Background
Language sampling has a long-standing history in linguistic typology. Rijkhoff et al. (1993) state that testing a linguistic hypothesis on a few similar languages is not sufficient for drawing conclusions on human language as a whole. Rather, they argue that generalizable samples should capture the diversity of the world’s languages. Language samples can be constructed in a number of ways. Rijkhoff and Bakker (1998) distinguish three types of sampling: random, probability and variety. Random samples do not have any specific stratification criteria, whereas probability samples are constructed to minimize bias, for instance by avoiding a skew towards languages from particular language families or geographical areas. Lastly, variety samples are designed to contain as much linguistic variety as possible, including relatively uncommon typological characteristics (Miestamo et al., 2016).
Language sampling methods in NLP have a similar goal: to sample languages in such a way that the findings (e.g., multilingual performance estimates) are expected to generalize across languages. Yet, their justifications are commonly less rigorous than those in linguistic typology. For language sampling methods in NLP, it is usually not explicitly mentioned which type of sampling is conducted, and for which reasons. In practice, most papers do not provide a justification for their selections (see §4.3), which resembles random sampling. A rare exception to this is Ponti et al. (2020), who use variety sampling and mention that this ensures evaluation that is robust towards typological characteristics that were likely not seen during training.
@@
While the goal of language sampling for ensuring generalizable conclusions is similar between NLP and linguistic typology, their resources are typically different. In linguistic typology, the variables under investigation are often typological characteristics. A typologist may aim to test whether a correlation of features holds across languages. Such a correlation could then be referred to as a ‘universal’: languages with object-verb ordering would be postpositional Greenberg (1963), for instance. To avoid circularity, language sampling is not done directly based on those features. Instead, popular methods rely on (combinations of) language family, genus, and geography (Miestamo et al., 2016).
By contrast, what is evaluated in multilingual NLP does not concern typological features directly, but rather, model performance across languages. In fact, to support generalizability claims, multilingual benchmarks must explicitly contain a range of typological phenomena (Bender, 2009, 2011; Pikuliak and Simko, 2022). Relying on geographical and genetic relations between languages is not guaranteed to ensure this. Nichols (1996) argues that “the kind of variables that define genealogical groups and tree shapes have a very different nature from the kind of variables that define typological diversity.” For estimating the generalizability of a language sample in NLP evaluation, it is possible, and even desired, to have direct insight into the actual typological properties that are included in the selection (see also §6).
@@ 3.2 Annotation Guidelines
    Which languages are selected?
@@ 3.3 Inter-Annotator Agreement
For the annotations concerning claims (1) and datasets (2), we achieve substantial agreement. We resolve the disagreements by discussing the annotations and merge them into a single label. For the sake of transparency, and to support incentives for human label variation Plank (2022), we also release the non-aggregated annotations. [^4] At the repository listed below the abstract. The agreement for the language selection annotation (3) is somewhat low, which is partially explained by the fact that we calculate agreement over all languages together, rather than on the micro-level for individual languages. This is because there is no ground truth regarding the number of languages to annotate. Furthermore, a number of inconsistencies are due to ISO-639-3 variants (jap $\rightarrow$ jpn or ger $\rightarrow$ deu) and different codes for possibly ambiguous language mentions (‘Norwegian’ could be nor or nob). We iteratively resolve these issues to get our final list of languages per paper.
@@ 4.1 Publication Overview
Figure 2: Number of papers with a claim by venue.
@@ 5 Typological Analysis
Figure 5: Distributions of mean pairwise lang2vec distances and feature inclusion per paper. On the left are approximations based on common justifications for claiming ‘typological diversity’: geography and genealogy. On the right two different approximations based on typological features: MPSD and Grambank feature value inclusion.
From Section 4.2), it follows that there is a geographical skew, where languages from certain areas are over-represented (Figure 4). This is not surprising, as it is often more feasible to select languages for which there are existing resources and typological descriptions, than languages for which these have to be gathered from scratch. This constitutes bibliographic bias (Rijkhoff and Bakker, 1998). However, we cannot draw conclusions about typological diversity on this basis, as typology is not equal to geography, nor phylogeny (Cysouw, 2013). Therefore, in this section, we analyze the typological diversity according to metrics that take into account syntactic features, and the absolute included typological feature values. Ponti et al. (2020) previously quantified language diversity with typology (URIEL), geography and family indices. These metrics do not take into account the distance between pairs of languages, and thus do not provide insights into the independence of languages in a sample. Additionally, no code or clear descriptions are provided, which limits their use.
@@ 5.1 Mean Language Distance
The typological distances vary considerably, with outliers on either side (Figure 5). In papers claiming typological diversity, the minimum MPSD is found in Goel et al. (2022), who use English, French and Spanish. The maximum MPSD is found for Vania et al. (2019) their selection of North Sámi, Galician, and Kazah.
@@ Acknowledgements
We thank the TypNLP group, especially Heather Lent, for the insightful feedback on an early version of this paper. We also thank Thomas Bauwens for his help in creating the tables (using fiject).

### v4

@@ 1 Introduction
Current work increasingly evaluates models on multiple languages, but because of resource constraints, it is not realistic to test a model on the thousands of languages in the world. In order to still ensure a degree of generalizability, previous work recognizes the importance of diverse language sampling. Ponti et al. (2020) suggest that merely evaluating on a small set of similar languages is an unreliable method for estimating a multilingual model’s performance, since such evaluations lack robustness to previously unseen typological properties, for instance. In an attempt to address these sampling issues, NLP research increasingly claims to rely on linguistic typology (see Figure 1). Evaluating on a set of so-called ‘typologically diverse’ languages implies robustness over typological variations, implicitly claiming broad generalizability across languages. Importantly, however, there is a fundamental lack of common understanding of exactly what a claim of typological diversity entails.
In this work, we systematically investigate how the term ‘typological diversity’ is used in NLP research by surveying the literature for past and current practices. We find that there is no set definition of ‘typological diversity’ in NLP: justifications for claims range from (i) none whatsoever, to (ii) using language phylogeny as a proxy for typology, to (iii) basing diversity on a more or less well-founded subset of available typological features. As a consequence, typological language similarity and typological feature value inclusion vary considerably across papers (§5). Furthermore, we demonstrate why this is problematic, by showing how a skewed selection of evaluation languages can lead to overstated multilingual performance and poor generalizability (§6). We recommend future work using the term ‘typologically diverse’ to document their language samples, and to add a measure of typological diversity, such as the mean pairwise language distance (§5.1) or typological feature inclusion (§5.2). Better-documented language sampling is crucial for the development of multilingual NLP, because it provides insights into commonly overlooked linguistic difficulties of multilingual language modelling. We provide software to verify and visualize this, linked below the abstract.
@@ 2 Background
Language sampling has a long-standing history in linguistic typology. Rijkhoff et al. (1993) state that testing a linguistic hypothesis on a few similar languages is not sufficient for drawing conclusions on human language as a whole. Rather, they argue that generalizable samples should capture the diversity of the world’s languages. Language samples can be constructed in a number of ways. Rijkhoff and Bakker (1998) distinguish three types of sampling: random, probability and variety. Random samples do not have any specific stratification criteria, whereas probability samples are constructed to minimize bias, for instance by avoiding a skew towards languages from particular language families or geographical areas. Variety samples are designed to contain as much linguistic variety as possible, including relatively uncommon typological characteristics (Miestamo et al., 2016). Finally, in addition to the three methods outlined, sampling can be based on data availability, which is defined as convenience sampling (Velupillai, 2012, p. 50).
Language sampling methods in NLP have a similar goal: to sample languages in such a way that the findings (e.g., multilingual performance estimates) are expected to generalize across languages. Yet, their justifications are commonly less rigorous than those in linguistic typology. For language sampling methods in NLP, it is usually not explicitly mentioned which type of sampling is conducted, and for which reasons. In practice, most papers do not provide a justification for their selections (see §4.3). This is likely caused by data availability, albeit implicitly, which resembles convenience sampling. A rare exception to this is Ponti et al. (2020), who use variety sampling and mention that this ensures evaluation that is robust towards typological characteristics that were likely not seen during training.
@@
While the goal of language sampling for ensuring generalizable conclusions is similar between NLP and linguistic typology, their resources are typically different. In linguistic typology, the variables under investigation are often typological characteristics. A typologist may aim to test whether a correlation of features holds across languages. Such a correlation could then be referred to as a ‘universal’: languages with object-verb ordering would be postpositional Greenberg (1963), for instance. To avoid circularity, language sampling is not done directly based on those features. Instead, methods commonly rely on (combinations of) language family, genus, and geography (Miestamo et al., 2016).
By contrast, what is evaluated in multilingual NLP does not concern typological features directly, rather, model performance across languages. In fact, to support claims of generalizability, multilingual benchmarks must explicitly contain a range of typological phenomena (Bender, 2009, 2011; Pikuliak and Simko, 2022). Relying on geographical and genetic relations between languages is not guaranteed to ensure this. Stoll and Bickel (2013) argue that “the kind of variables that define genealogical groups and tree shapes have a very different nature from the kind of variables that define typological diversity”, based on work from Nichols (1996). For estimating the generalizability of a language sample in NLP evaluation, it is possible, and even desired, to have direct insight into the actual typological properties that are included in the selection (see also §6).
This motivation is similarly voiced by Samardzic et al. (2024), who introduce a data-driven method of relating language samples in particular datasets to a reference language sample that is assumed to be typologically diverse. Yet, our work differs in crucial ways. Our aims are to look at what it means to make claims of typological diversity in current research practices and how to characterize the language samples such claims are made about. Our methods work as standalone measures; we do not need a reference sample to compare against. In addition, our approach is centered around languages, thus only requiring language descriptions, not entire datasets. Samardzic et al. (2024) approximate ‘aggregated morphological features’ through the average word length in a corpus, but this number may vary across domains within the same language.
@@ 3.2 Annotation Guidelines
    Which languages are included?
@@ 3.3 Inter-Annotator Agreement
For the annotations concerning claims (1) and datasets (2), we achieve substantial agreement. We resolve the disagreements by discussing the annotations and merging them into a single label. For the sake of transparency, and to support incentives for human label variation Plank (2022), we also release the non-aggregated annotations. [^4] At the repository listed below the abstract. The agreement for the language selection annotation (3) is somewhat low, which is partially explained by the fact that we calculate agreement over all languages together, rather than on the micro-level for individual languages. This is because there is no ground truth regarding the number of languages to annotate. Furthermore, a number of inconsistencies are due to ISO-639-3 variants (jap $\rightarrow$ jpn or ger $\rightarrow$ deu) and different codes for possibly ambiguous language mentions (‘Norwegian’ could be nor or nob). We iteratively resolve these issues to get our final list of languages per paper.
Figure 2: Number of papers with a claim by venue.
@@ 5 Typological Analysis
Figure 5: Distributions of mean pairwise lang2vec distances and feature inclusion per paper. On the left are approximations based on common justifications for claiming ‘typological diversity’: geography ($\mu=0.28$, $\sigma=0.11$) and genealogy ($\mu=0.94$, $\sigma=0.05$). On the right two different approximations based on typological features: MPSD ($\mu=0.64$, $\sigma=0.07$) and Grambank feature value inclusion ($\mu=0.72$, $\sigma=0.17$).
From Section 4.2, it follows that there is a geographical skew, where languages from certain areas are over-represented (Figure 4). This is not surprising, as it is often more feasible to select languages for which there are existing resources and typological descriptions, than languages for which these have to be gathered from scratch. This constitutes bibliographic bias (Rijkhoff and Bakker, 1998). However, we cannot draw conclusions about typological diversity on this basis, as typology is not equal to geography, nor phylogeny (Cysouw, 2013). Therefore, in this section, we analyze the typological diversity according to metrics that take into account syntactic features, and the absolute included typological feature values. Ponti et al. (2020) previously quantified language diversity with typology (URIEL), geography, and family indices. These metrics do not take into account the distance between pairs of languages, and thus do not provide insights into the independence of languages in a sample. Additionally, no code or clear descriptions are provided, which limits their use.
@@ 5.1 Mean Language Distance
The typological distances vary considerably, with outliers on either side (Figure 5). In papers claiming typological diversity, the minimum MPSD is found in Goel et al. (2022), who use English, French, and Spanish. The maximum MPSD is found for Vania et al. (2019) their selection of North Sámi, Galician, and Kazah.
@@ Acknowledgements
We thank the TypNLP group at Aalborg University and the LAGoM-NLP group at KU Leuven for insightful feedback on earlier versions of this paper, in particular Heather Lent. We also thank Thomas Bauwens for his help in creating the tables (using fiject).